- Current status as of 1/22: IR and Green reached X end PD.
1/22: Surface inspection for PR mirrors. Cleaning will be done after 1/31.
1/23: Surface inspection for SR mirrors.
1/24-25: Cleaning HR of SR3, SRM.
1/23: BS is aligned to Y-end PD by IR.
1/27: ITMY is aligned to REFL by IR.
1/27: ITMX is aligned to REFL by IR.
1/27: Beam position of IR at SR3 is confirmed how far from the center.
1/28-29: SR3 is aligned to SR2 center by IR.
1/28: Alignment of SR2 -> SRM by IR.
1/29: Green beam is injected from SR2 AR surface to Y end PD.
1/31: Removing bellows at PR and SR area.
1/31-2/1: Cleaning HR of PR3 and HR, AR of SR2
2/5: Alignment of SRM to OMMT1 by IR
2/6-8: OMMT -> OSTM by IR, OSTM -> OMC by IR
2/6-8: Alignment of ETMY -> POS by Green
- 2/12-15: Closing door for central area, pumping down
- needs drift monitor for VIS
- Today is the last day of Wang-san.
- Report klog, and gokagra ML if urgent things.
- If you work other location from where it was confirmed at the morning meeting, please contact related poeple.
Report of work
- Type B: Height of SRs were adjusted.
- COM: Beam was aligned to OMMT1 but had 2-3cm off from center.
- OMMT1 suspension will be moved on Friday.
- Beam reached to OMMT2 already but off from the center in 2-3cm as the same oder as OMMT1.
- ETMY: Mechanical trouble or something caused that ETMY did not work well. TF was something wrong.
- Oplev beam of SR? was hit at edge of optical table but it was fixed.
- SR models was splited.
- Pcal: QPD at ETMY was installed.
- COM: ETMY alignment by green.
- PCal: Needs little more ealignment.